# Inversion of adjunction for quotient singularities III: semi-invariant case

### 10 Dec 2023 v1

@@ 3. Settings and Denef and Loeser’s theory
    We say that $f$ is $G$-semi-invariant if for any $\alpha\in G$, there exists $10\leq a\leq d-$ 1 such that $\alpha(f)=\xi^{a}f$.
@@
    Suppose that $f$ is $G$-semi-invariant. For $\alpha\in G$, we define $w_{\alpha}(f):=\frac{a}{d}$ when $\alpha(f)=\xi^{a}f$ with $10\leq a\leq d-$ 1.
Let $c$ be a non-negative integer with $c\leq N$. We set $n:=N-c$. Let $f_{1},\ldots,f_{c}\in R$ be a regular sequence which is contained in the maximal ideal $(x_{1},\ldots,x_{N})$ at the origin. Suppose that $f_{1},\ldots,f_{c}$ are $G$-semi-invariant. We set
@@
Let $\gamma\in G$. Since $G$ is a finite group, $\gamma$ can be diagonalized with some new basis $x_{1}^{(\gamma)},\ldots,x_{N}^{(\gamma)}$ those are $k$-linear combinations of original basis $x_{1},\ldots,x_{N}$. Let $\operatorname{diag}\left(\xi^{e_{1}},\ldots,\xi^{e_{N}}\right)$ be the diagonal matrix with $10\leq e_{i}\leq d-$ 1. Then, the age $\operatorname{age}(\gamma)$ of $\gamma$ is defined by $\operatorname{age}(\gamma):=\frac{1}{d}\sum_{i=1}^{N}e_{i}$. We also define a $k$-algebra homomorphism ${\lambda}^{*}_{\gamma}:R\to R[t^{1/d}]$ as follows:
@@
We note that the ring homomorphism ${\lambda}^{*}_{\gamma}$ does not depend on the choice of the basis $x_{1}^{(\gamma)},\ldots,x_{N}^{(\gamma)}$ as long as $\gamma$ is diagonalized by this basis. Let $C_{\gamma}$ denote the centralizer of $\gamma$ in $G$.
@@
[…]
@@
    If $f\in R$ is a $G$-semi-invariant element, then for any $\alpha\in G$, we have $\lambda^{*}_{\alpha}(f)t^{-w_{\alpha}(f)}\in R[t]$.
@@
Let $\alpha\in C_{\gamma}$. Then, by replacing the bases $x_{1}^{(\gamma)},\ldots,x_{N}^{(\gamma)}$, we may assume that $\alpha$ is also a diagonal matrix with this basis. Then, the assertion (1) is obvious.
@@
$\left(t\cdot\lambda^{*}_{\gamma}(f_{i})t^{-w_{\gamma}(f_{i})}\right)^{d}=t^{d-dw_{\gamma}(f_{i})}\lambda^{*}_{\gamma}(f_{i}^{d})\in\lambda^{*}_{\gamma}(I_{B})R[t],$
@@
By Lemmas 3.4(2) and 3.5(1), we have the following diagram of rings:
[…]
@@
$\widetilde{B}^{(\gamma)}$ $:=\operatorname{Spec}\left(R^{C_{\gamma}}[t]/\left(I_{\overline{B}^{(\gamma)}}\cap R^{C_{\gamma}}[t]\right)\right),\quad$ $\overline{B}^{(\gamma)}$ $:=\operatorname{Spec}\left(R[t]/I_{\overline{B}^{(\gamma)}}\right).$
@@
We identify $B_{\infty}$, $\widetilde{B}_{\infty}^{(\gamma)}$ and $\overline{B}_{\infty}^{(\gamma)}$ with the closed subset of $A_{\infty}$, $\widetilde{A}^{(\gamma)}_{\infty}$ and $\overline{A}^{(\gamma)}_{\infty}$ (cf. [NS22]*Lemma 2.28(1)).
@@ 4. Order calculations
[…]
@@
[…]
@@
[…]
Here, we defined $k[t^{1/d}]$-schemes $A_{i}$ for $i=1,2,3$ and $B_{i}$ for $i=1,2,3,4$ as follows:
$A_{1}$ $:=\operatorname{Spec}\left(R^{G}[t^{1/d}]\right),$ $A_{2}:=A_{3}:=\operatorname{Spec}\left(R[t^{1/d}]\right),$
$B_{1}$ $:=\operatorname{Spec}\left(R^{G}[t^{1/d}]\big{/}\left(I_{B}R^{G}[t^{1/d}]\right)\right),$ $B_{2}:=\operatorname{Spec}\left(R[t^{1/d}]\big{/}\left(I_{\overline{B}}R[t^{1/d}]\right)\right),$
$B_{3}$ $:=\operatorname{Spec}\left(R[t^{1/d}]\big{/}\left(\lambda^{*}_{\gamma}\left(I_{\overline{B}}\right)R[t^{1/d}]\right)\right),$ $B_{4}:=\operatorname{Spec}\left(R[t^{1/d}]\big{/}\left(I_{\overline{B}^{(\gamma)}}R[t^{1/d}]\right)\right).$
@@
$B_{4}=\overline{B}^{(\gamma)}\times_{\operatorname{Spec}k[t]}\operatorname{Spec}k[t^{1/d}].$
@@
By definition, the morphisms $B_{4}\to B_{3}$ and $B_{3}\to B_{2}$ are isomorphic outside $t^{1/d}=0$. Then, the étale morphism $\overline{B}\times\left(\mathbb{A}^{1}_{k}\setminus\{0\}\right)\to\overline{B}^{(\gamma)}_{t\not=0}$ in Remark 3.6(3) is given by the compositions of the following mophisms:
@@
$\overline{B}\simeq\overline{B}\times\{b\}\xrightarrow{\ \ \simeq\ \ }\overline{B}^{(\gamma)}_{t=a}:=\operatorname{Spec}\left(R[t]/\left(I_{\overline{B}^{(\gamma)}}+(t-a)\right)\right).$
@@
We define a sheaf $L_{4}$ on $B_{4}$ by
$L_{4}:=t^{w_{\gamma}({\bf f})}\left(\operatorname{det}^{-1}\left(I_{2}/I_{2}^{2}\right)\right)\big{|}_{B_{4}}\otimes_{\mathcal{O}_{B_{4}}}\bigl{(}\Omega_{A_{3}/k[t^{1/d}]}^{N}\bigr{)}\big{|}_{B_{4}},$
$\text{where \quad}I_{2}:=I_{\overline{B}}R[t^{1/d}],\quad w_{\gamma}({\bf f}):=w_{\gamma}(f_{1})+\cdots+w_{\gamma}(f_{c}).$
Here, $\operatorname{det}^{-1}\left(I_{2}/I_{2}^{2}\right)=\mathcal{H}om_{\mathcal{O}_{B_{2}}}\left(\bigwedge^{c}(I_{2}/I_{2}^{2}),\mathcal{O}_{B_{2}}\right)$.
@@
    There is a natural morphism $\Omega^{n}_{B_{4}/k[t^{1/d}]}\to L_{4}$, and we have
    $\operatorname{Im}\left(\Omega^{n}_{B_{4}/k[t^{1/d}]}\to L_{4}\right)=\operatorname{Fitt}^{n}\left(\Omega_{B_{4}/k[t^{1/d}]}\right)\otimes_{\mathcal{O}_{B_{4}}}L_{4}.$
@@
    There is a natural morphism $t^{r\cdot w_{\gamma}({\bf f})}\left(\omega^{[r]}_{B_{1}/k[t^{1/d}]}\big{|}_{B_{4}}\right)\to L^{\otimes r}_{4}$, and we have
    $\operatorname{Im}\left(t^{r\cdot w_{\gamma}({\bf f})}\left(\omega^{[r]}_{B_{1}/k[t^{1/d}]}\big{|}_{B_{4}}\right)\to L^{\otimes r}_{4}\right)=t^{r\cdot\operatorname{age}(\gamma)}L^{\otimes r}_{4}.$
@@
We define an invertible sheaf $L_{3}$ on $B_{3}$ by
$L_{3}:=\left(\operatorname{det}^{-1}\left(I_{2}/I_{2}^{2}\right)\right)\big{|}_{B_{3}}\otimes_{\mathcal{O}_{B_{3}}}\left(\Omega_{A_{3}/k[t^{1/d}]}^{N}\right)\big{|}_{B_{3}}.$
Then, by the definitions of $L_{3}$ and $L_{4}$, we have a natural injective morphism $L_{4}\hookrightarrow L_{3}|_{B_{4}}$, and we have
$\operatorname{Im}\left(L_{4}\hookrightarrow L_{3}|_{B_{4}}\right)=t^{w_{\gamma}({\bf f})}\left(L_{3}|_{B_{4}}\right).$
@@
$\omega^{[r]}_{B_{1}/k[t^{1/d}]}\big{|}_{B_{2}}\simeq\omega^{\otimes r}_{B_{2}/k[t^{1/d}]}\simeq\left(\operatorname{det}^{-1}\left(I_{2}/I_{2}^{2}\right)\right)^{\otimes r}\otimes_{\mathcal{O}_{B_{2}}}\left(\Omega_{A_{2}/k[t^{1/d}]}^{N}\right)^{\otimes r}\big{|}_{B_{2}}.$
Furthermore, we have a natural morphism $\Omega_{A_{2}/k[t^{1/d}]}^{N}\big{|}_{A_{3}}\to\Omega_{A_{3}/k[t^{1/d}]}^{N}$, and its image is equal to $t^{\operatorname{age}(\gamma)}\left(\Omega_{A_{3}/k[t^{1/d}]}^{N}\right)$. Hence, we have a natural morphism $\omega^{[r]}_{B_{1}/k[t^{1/d}]}\big{|}_{B_{3}}\to L^{\otimes r}_{3}$, and its image is equal to
$\operatorname{Im}\left(\omega^{[r]}_{B_{1}/k[t^{1/d}]}\big{|}_{B_{3}}\to L^{\otimes r}_{3}\right)=t^{r\cdot\operatorname{age}(\gamma)}L^{\otimes r}_{3}.$
@@
$\Bigl{(}\bigwedge^{c}(I_{2}/I_{2}^{2})\Bigr{)}\big{|}_{B_{4}}\otimes_{\mathcal{O}_{B_{4}}}\Omega^{n}_{B_{4}/k[t^{1/d}]}\to\left(\Omega_{A_{3}/k[t^{1/d}]}^{N}\right)\big{|}_{B_{4}},$
@@
$t^{w_{\gamma}({\bf f})}\cdot\operatorname{Fitt}^{n}\left(\Omega_{B_{4}/k[t^{1/d}]}\right)\otimes\left(\Omega_{A_{3}/k[t^{1/d}]}^{N}\right)\big{|}_{B_{4}}.$
@@
$S_{2}:=R[t^{1/d}]\big{/}\left(I_{\overline{B}}R[t^{1/d}]\right),\quad S_{4}:=R[t^{1/d}]\big{/}\left(I_{\overline{B}^{(\gamma)}}R[t^{1/d}]\right),\quad I_{4}:=I_{\overline{B}^{(\gamma)}}R[t^{1/d}].$
Let $M\subset\Omega_{R[t^{1/d}]/k[t^{1/d}]}$ be the $R[t^{1/d}]$-submodule generated by $df$ for $f\in I_{4}$. Then, we have
$\left(\Omega_{R[t^{1/d}]/k[t^{1/d}]}\big{/}M\right)\otimes_{R[t^{1/d}]}S_{4}\simeq\Omega_{S_{4}/k[t^{1/d}]}.$
Since $I_{4}$ is generated by $c$ elements, we have $\bigwedge^{c+1}M=0$. Therefore, a natural map $\Omega_{R[t^{1/d}]/k[t^{1/d}]}^{c}\otimes\Omega_{R[t^{1/d}]/k[t^{1/d}]}^{N-c}\to\Omega_{R[t^{1/d}]/k[t^{1/d}]}^{N}$ induces a map
@@
$\left(I_{2}/I_{2}^{2}\right)\otimes_{S_{2}}S_{4}\to I_{4}/I_{4}^{2}\xrightarrow{\overline{h}\mapsto d(h)\otimes 1}M\otimes_{R[t^{1/d}]}S_{4}.$

### 30 Apr 2026 v2

@@ 3. Settings and Denef and Loeser’s theory
    We say that $f$ is $G$-semi-invariant if for any $\gamma\in G$, there exists $0\leq a\leq d-1$ such that $\gamma(f)=\xi^{a}f$.
@@
    Suppose that $f$ is $G$-semi-invariant. For $\gamma\in G$, we define $w_{\gamma}(f):=\frac{a}{d}$ when $\gamma(f)=\xi^{a}f$ with $0\leq a\leq d-1$.
Let $c$ be a non-negative integer with $c\leq N$. We set n:=Ncn:=N-c. Let $f_{1},\ldots,f_{c}\in R$ be a regular sequence which is contained in the maximal ideal $(x_{1},\ldots,x_{N})$ at the origin. Suppose that $f_{1},\ldots,f_{c}$ are $G$-semi-invariant. We set
@@
Let $\gamma\in G$. Since $G$ is a finite group, $\gamma$ can be diagonalized with respect to a new basis $x_{1}^{(\gamma)},\ldots,x_{N}^{(\gamma)}$ consisting of $k$-linear combinations of the original basis $x_{1},\ldots,x_{N}$. Let $\operatorname{diag}\left(\xi^{e_{1}},\ldots,\xi^{e_{N}}\right)$ be the diagonal matrix with $0\leq e_{i}\leq d-1$. Then, the age $\operatorname{age}(\gamma)$ of $\gamma$ is defined by $\operatorname{age}(\gamma):=\frac{1}{d}\sum_{i=1}^{N}e_{i}$. We also define a $k$-algebra homomorphism ${\lambda}^{*}_{\gamma}:R\to R[t^{1/d}]$ as follows:
@@
We note that the ring homomorphism ${\lambda}^{*}_{\gamma}$ does not depend on the choice of the basis $x_{1}^{(\gamma)},\ldots,x_{N}^{(\gamma)}$ as long as $\gamma$ is diagonalized with respect to this basis. Let $C_{\gamma}$ denote the centralizer of $\gamma$ in $G$.
Remark 3.4.
In [NS22]*Section 3, $\lambda^{*}_{\gamma}$ is defined as a $k[t]$-ring homomorphism from $R[t]^{G}$ to $R[t]^{C_{\gamma}}$ satisfying $\lambda^{*}_{\gamma}(x_{i}^{(\gamma)})=t^{e_{i}/d}x_{i}^{(\gamma)}$. Although the domain and codomain are different, we use the same symbol because the map in [NS22] is the restriction of the base change of our $\lambda^{*}_{\gamma}$ to $R[t]^{G}$.
@@
    For any $\delta\in C_{\gamma}$, the composition $R\xrightarrow{\delta}R\xrightarrow{\lambda^{*}_{\gamma}}R[t^{1/d}]$ coincides with the composition $R\xrightarrow{\lambda^{*}_{\gamma}}R[t^{1/d}]\xrightarrow{\delta}R[t^{1/d}]$. Here, $R[t^{1/d}]\xrightarrow{\delta}R[t^{1/d}]$ is the $k[t^{1/d}]$-ring homomorphism induced by $R\xrightarrow{\delta}R$.
@@
    If $f\in R$ is a $G$-semi-invariant element, then for any $\delta\in G$, we have $\lambda^{*}_{\delta}(f)t^{-w_{\delta}(f)}\in R[t]$.
@@
Let $\delta\in C_{\gamma}$. Then, by replacing the bases $x_{1}^{(\gamma)},\ldots,x_{N}^{(\gamma)}$, we may assume that $\delta$ is also a diagonal matrix with respect to this basis. Then, the assertion (1) is obvious.
@@
$\bigl(t\cdot\lambda^{*}_{\gamma}(f_{i})t^{-w_{\gamma}(f_{i})}\bigr)^{d}=t^{d-dw_{\gamma}(f_{i})}\lambda^{*}_{\gamma}(f_{i}^{d})\in\lambda^{*}_{\gamma}(I_{B})R[t],$
@@
By Lemmas 3.5(2) and 3.6(1), we have the following diagram of rings:
${R^{G}[t]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{x_{i}^{(\gamma)}\mapsto t^{e_{i}/d}x_{i}^{(\gamma)}}$${R^{C_{\gamma}}[t]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R[t]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R^{G}[t]/\left(I_{B}R^{G}[t]\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R^{C_{\gamma}}[t]/\left(I_{\overline{B}^{(\gamma)}}\cap R^{C_{\gamma}}[t]\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R[t]/I_{\overline{B}^{(\gamma)}}}$
@@
$\widetilde{B}^{(\gamma)}$ $:=\operatorname{Spec}\left(R^{C_{\gamma}}[t]/\bigl(I_{\overline{B}^{(\gamma)}}\cap R^{C_{\gamma}}[t]\bigr)\right),\quad$ $\overline{B}^{(\gamma)}$ $:=\operatorname{Spec}\bigl(R[t]/I_{\overline{B}^{(\gamma)}}\bigr).$
@@
We identify $B_{\infty}$, $\widetilde{B}_{\infty}^{(\gamma)}$ and $\overline{B}_{\infty}^{(\gamma)}$ with the corresponding closed subset of $A_{\infty}$, $\widetilde{A}^{(\gamma)}_{\infty}$ and $\overline{A}^{(\gamma)}_{\infty}$ (cf. [NS22]*Lemma 2.28(1)).
@@ 4. Order calculations
${R^{G}[t^{1/d}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R[t^{1/d}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{x_{i}^{(\gamma)}\mapsto t^{e_{i}/d}x_{i}^{(\gamma)}}$${R[t^{1/d}]}$${I_{B}R^{G}[t^{1/d}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\bigcup}$${I_{\overline{B}}R[t^{1/d}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\bigcup}$${I_{\overline{B}^{(\gamma)}}R[t^{1/d}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\bigcup}$${\left(\lambda^{*}_{\gamma}(f_{1}),\ldots,\lambda^{*}_{\gamma}(f_{c})\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\bigcup}$
@@
${R^{G}[t^{1/d}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R[t^{1/d}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{x_{i}^{(\gamma)}\mapsto t^{e_{i}/d}x_{i}^{(\gamma)}}$${R[t^{1/d}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R^{G}[t^{1/d}]/\left(I_{B}R^{G}[t^{1/d}]\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R[t^{1/d}]/\left(I_{\overline{B}}R[t^{1/d}]\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R[t^{1/d}]/\left(\lambda^{*}_{\gamma}\left(I_{\overline{B}}\right)R[t^{1/d}]\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R[t^{1/d}]\big/\left(I_{\overline{B}^{(\gamma)}}R[t^{1/d}]\right)}$
@@
${A_{k[t^{1/d}]}=A\times\mathbb{A}^{1}_{k}}$${\overline{A}_{k[t^{1/d}]}=\overline{A}\times\mathbb{A}^{1}_{k}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\overline{A}^{(\gamma)}_{1/d}=\mathbb{A}^{N}_{k}\times\mathbb{A}^{1}_{k}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${B_{k[t^{1/d}]}=B\times\mathbb{A}^{1}_{k}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\overline{B}_{k[t^{1/d}]}=\overline{B}\times\mathbb{A}^{1}_{k}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widehat{B}^{(\gamma)}_{1/d}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\overline{B}^{(\gamma)}_{1/d}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
Here, we considered the following $k[t]$-schemes:
$A_{k[t^{1/d}]}$ $:=\operatorname{Spec}\bigl(R^{G}[t^{1/d}]\bigr),$ $\overline{A}_{k[t^{1/d}]}:=\overline{A}^{(\gamma)}_{1/d}:=\operatorname{Spec}\bigl(R[t^{1/d}]\bigr),$
$B_{k[t^{1/d}]}$ $:=\operatorname{Spec}\left(R^{G}[t^{1/d}]\big/\bigl(I_{B}R^{G}[t^{1/d}]\bigr)\right),$ $\overline{B}_{k[t^{1/d}]}:=\operatorname{Spec}\left(R[t^{1/d}]\big/\bigl(I_{\overline{B}}R[t^{1/d}]\bigr)\right),$
$\widehat{B}^{(\gamma)}_{1/d}$ $:=\operatorname{Spec}\left(R[t^{1/d}]\big/\bigl(\lambda^{*}_{\gamma}\left(I_{\overline{B}}\right)R[t^{1/d}]\bigr)\right),$ $\overline{B}^{(\gamma)}_{1/d}:=\operatorname{Spec}\left(R[t^{1/d}]\big/\bigl(I_{\overline{B}^{(\gamma)}}R[t^{1/d}]\bigr)\right).$
@@
$\overline{B}^{(\gamma)}_{1/d}=\overline{B}^{(\gamma)}\times_{\operatorname{Spec}k[t]}\operatorname{Spec}k[t^{1/d}].$
@@
By definition, the morphisms $\overline{B}^{(\gamma)}_{1/d}\to\widehat{B}^{(\gamma)}_{1/d}$ and $\widehat{B}^{(\gamma)}_{1/d}\to\overline{B}_{k[t^{1/d}]}$ are isomorphic outside $t^{1/d}=0$. Then, the étale morphism $\overline{B}\times\left(\mathbb{A}^{1}_{k}\setminus\{0\}\right)\to\overline{B}^{(\gamma)}_{t\not=0}$ in Remark 3.7(3) is given by the compositions of the following mophisms:
@@
$\overline{B}\simeq\overline{B}\times\{b\}\xrightarrow{\ \ \simeq\ \ }\overline{B}^{(\gamma)}_{t=a}:=\operatorname{Spec}\left(R[t]/\bigl(I_{\overline{B}^{(\gamma)}}+(t-a)\bigr)\right).$
@@
We define a sheaf $L_{\overline{B}^{(\gamma)}_{1/d}}$ on $\overline{B}^{(\gamma)}_{1/d}$ by
$L_{\overline{B}^{(\gamma)}_{1/d}}:=t^{w_{\gamma}({\bf f})}\left(\operatorname{det}^{-1}\Bigl(I_{\overline{B}_{k[t^{1/d}]}}/I^{2}_{\overline{B}_{k[t^{1/d}]}}\Bigr)\right)\big|_{\overline{B}^{(\gamma)}_{1/d}}\otimes_{\mathcal{O}_{\overline{B}^{(\gamma)}_{1/d}}}\Bigl(\Omega_{\overline{A}^{(\gamma)}_{1/d}/k[t^{1/d}]}^{N}\Bigr)\big|_{\overline{B}^{(\gamma)}_{1/d}}$,
$\text{where \quad}I_{\overline{B}_{k[t^{1/d}]}}:=I_{\overline{B}}R[t^{1/d}],\quad w_{\gamma}({\bf f}):=w_{\gamma}(f_{1})+\cdots+w_{\gamma}(f_{c}).$
Here, $\operatorname{det}^{-1}\Bigl(I_{\overline{B}_{k[t^{1/d}]}}/I^{2}_{\overline{B}_{k[t^{1/d}]}}\Bigr)=\mathcal{H}om_{\mathcal{O}_{\overline{B}_{k[t^{1/d}]}}}\left(\bigwedge^{c}\Bigl(I_{\overline{B}_{k[t^{1/d}]}}/I^{2}_{\overline{B}_{k[t^{1/d}]}}\Bigr),\mathcal{O}_{\overline{B}_{k[t^{1/d}]}}\right)$.
Recall that $r$ is a positive integer such that $\omega_{B}^{[r]}$ is invertible.
@@
    There is a natural morphism $\Omega^{n}_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\to L_{\overline{B}^{(\gamma)}_{1/d}}$, and we have
    $\operatorname{Im}\left(\Omega^{n}_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\to L_{\overline{B}^{(\gamma)}_{1/d}}\right)=\operatorname{Fitt}^{n}\Bigl(\Omega_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\Bigr)\otimes_{\mathcal{O}_{\overline{B}^{(\gamma)}_{1/d}}}L_{\overline{B}^{(\gamma)}_{1/d}}$.
@@
    There is a natural morphism $t^{r\cdot w_{\gamma}({\bf f})}\Bigl(\omega^{[r]}_{B_{k[t^{1/d}]}/k[t^{1/d}]}\big|_{\overline{B}^{(\gamma)}_{1/d}}\Bigr)\to L_{\overline{B}^{(\gamma)}_{1/d}}^{\otimes r}$, and we have
    $\operatorname{Im}\left(t^{r\cdot w_{\gamma}({\bf f})}\Bigl(\omega^{[r]}_{B_{k[t^{1/d}]}/k[t^{1/d}]}\big|_{\overline{B}^{(\gamma)}_{1/d}}\Bigr)\to L_{\overline{B}^{(\gamma)}_{1/d}}^{\otimes r}\right)=t^{r\cdot\operatorname{age}(\gamma)}L_{\overline{B}^{(\gamma)}_{1/d}}^{\otimes r}.$
@@
We define an invertible sheaf $L_{\widehat{B}^{(\gamma)}_{1/d}}$ on $\widehat{B}^{(\gamma)}_{1/d}$ by
$L_{\widehat{B}^{(\gamma)}_{1/d}}:=\left(\operatorname{det}^{-1}\Bigl(I_{\overline{B}_{k[t^{1/d}]}}/I^{2}_{\overline{B}_{k[t^{1/d}]}}\Bigr)\right)\big|_{\widehat{B}^{(\gamma)}_{1/d}}\otimes_{\mathcal{O}_{\widehat{B}^{(\gamma)}_{1/d}}}\left(\Omega_{\overline{A}^{(\gamma)}_{1/d}/k[t^{1/d}]}^{N}\right)\big|_{\widehat{B}^{(\gamma)}_{1/d}}$.
Then, by the definitions of $L_{\widehat{B}^{(\gamma)}_{1/d}}$ and $L_{\overline{B}^{(\gamma)}_{1/d}}$, we have a natural injective morphism $L_{\overline{B}^{(\gamma)}_{1/d}}\hookrightarrow L_{\widehat{B}^{(\gamma)}_{1/d}}|_{\overline{B}^{(\gamma)}_{1/d}}$, and we have
$\operatorname{Im}\left(L_{\overline{B}^{(\gamma)}_{1/d}}\hookrightarrow L_{\widehat{B}^{(\gamma)}_{1/d}}|_{\overline{B}^{(\gamma)}_{1/d}}\right)=t^{w_{\gamma}({\bf f})}\left(L_{\widehat{B}^{(\gamma)}_{1/d}}|_{\overline{B}^{(\gamma)}_{1/d}}\right)$.
@@
$\omega^{[r]}_{B_{k[t^{1/d}]}/k[t^{1/d}]}\big|_{\overline{B}_{k[t^{1/d}]}}$
$\simeq\omega^{\otimes r}_{\overline{B}_{k[t^{1/d}]}/k[t^{1/d}]}$
$\simeq\left(\operatorname{det}^{-1}\Bigl(I_{\overline{B}_{k[t^{1/d}]}}/I^{2}_{\overline{B}_{k[t^{1/d}]}}\Bigr)\right)^{\otimes r}\otimes_{\mathcal{O}_{\overline{B}_{k[t^{1/d}]}}}\Bigl(\Omega_{\overline{A}_{k[t^{1/d}]}/k[t^{1/d}]}^{N}\Bigr)^{\otimes r}\big|_{\overline{B}_{k[t^{1/d}]}}.$
Furthermore, we have a natural morphism $\Omega_{\overline{A}_{k[t^{1/d}]}/k[t^{1/d}]}^{N}\big|_{\overline{A}^{(\gamma)}_{1/d}}\to\Omega_{\overline{A}^{(\gamma)}_{1/d}/k[t^{1/d}]}^{N}$, and its image is equal to $t^{\operatorname{age}(\gamma)}\left(\Omega_{\overline{A}^{(\gamma)}_{1/d}/k[t^{1/d}]}^{N}\right)$. Hence, we have a natural morphism $\omega^{[r]}_{B_{k[t^{1/d}]}/k[t^{1/d}]}\big|_{\widehat{B}^{(\gamma)}_{1/d}}\to L_{\widehat{B}^{(\gamma)}_{1/d}}^{\otimes r}$, and its image is equal to
$\operatorname{Im}\left(\omega^{[r]}_{B_{k[t^{1/d}]}/k[t^{1/d}]}\big|_{\widehat{B}^{(\gamma)}_{1/d}}\to L_{\widehat{B}^{(\gamma)}_{1/d}}^{\otimes r}\right)=t^{r\cdot\operatorname{age}(\gamma)}L_{\widehat{B}^{(\gamma)}_{1/d}}^{\otimes r}.$
@@
$\left(\bigwedge^{c}\left(I_{\overline{B}_{k[t^{1/d}]}}/I^{2}_{\overline{B}_{k[t^{1/d}]}}\right)\right)\big|_{\overline{B}^{(\gamma)}_{1/d}}\otimes_{\mathcal{O}_{\overline{B}^{(\gamma)}_{1/d}}}\Omega^{n}_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\to\left(\Omega_{\overline{A}^{(\gamma)}_{1/d}/k[t^{1/d}]}^{N}\right)\big|_{\overline{B}^{(\gamma)}_{1/d}}$,
@@
$t^{w_{\gamma}({\bf f})}\cdot\operatorname{Fitt}^{n}\left(\Omega_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\right)\otimes\left(\Omega_{\overline{A}^{(\gamma)}_{1/d}/k[t^{1/d}]}^{N}\right)\big|_{\overline{B}^{(\gamma)}_{1/d}}$.
@@
$S_{\overline{B}_{k[t^{1/d}]}}$ $:=R[t^{1/d}]\big/\left(I_{\overline{B}}R[t^{1/d}]\right),$
$S_{\overline{B}^{(\gamma)}_{1/d}}$ $:=R[t^{1/d}]\big/\left(I_{\overline{B}^{(\gamma)}}R[t^{1/d}]\right),$
$I_{\overline{B}^{(\gamma)}_{1/d}}$ $:=I_{\overline{B}^{(\gamma)}}R[t^{1/d}].$
Let $M\subset\Omega_{R[t^{1/d}]/k[t^{1/d}]}$ be the $R[t^{1/d}]$-submodule generated by $df$ for $f\in I_{\overline{B}^{(\gamma)}_{1/d}}$. Then, we have
$\left(\Omega_{R[t^{1/d}]/k[t^{1/d}]}\big/M\right)\otimes_{R[t^{1/d}]}S_{\overline{B}^{(\gamma)}_{1/d}}\simeq\Omega_{S_{\overline{B}^{(\gamma)}_{1/d}}/k[t^{1/d}]}.$
Since $I_{\overline{B}^{(\gamma)}_{1/d}}$ is generated by $c$ elements, we have $\bigwedge^{c+1}M=0$. Therefore, a natural map $\Omega_{R[t^{1/d}]/k[t^{1/d}]}^{c}\otimes\Omega_{R[t^{1/d}]/k[t^{1/d}]}^{N-c}\to\Omega_{R[t^{1/d}]/k[t^{1/d}]}^{N}$ induces a map
@@
$\left(I_{\overline{B}_{k[t^{1/d}]}}/I^{2}_{\overline{B}_{k[t^{1/d}]}}\right)\otimes_{S_{\overline{B}_{k[t^{1/d}]}}}S_{\overline{B}^{(\gamma)}_{1/d}}\to I_{\overline{B}^{(\gamma)}_{1/d}}/I^{2}_{\overline{B}^{(\gamma)}_{1/d}}\xrightarrow{\overline{h}\mapsto d(h)\otimes 1}M\otimes_{R[t^{1/d}]}S_{\overline{B}^{(\gamma)}_{1/d}}$.